OpenStudy (anonymous):

Contsruct a graph after finding the following: A:Domain B:X and Y intercepts C:Symetry D:Asymptotes E:Interval of Increase/Decrease F:Local Max/Min G:Curvature and Inflection points

OpenStudy (anonymous):

\[y=\sin^3x\]

OpenStudy (anonymous):

Domain is all real numbers.

OpenStudy (anonymous):

x intercept =

OpenStudy (anonymous):

\[0=\sin^3x\]

OpenStudy (anonymous):

\[\pi(n)\]

OpenStudy (anonymous):

y intercept = 0

OpenStudy (anonymous):

symetry = odd about the orgin

OpenStudy (anonymous):

no asymptotes

OpenStudy (anonymous):

\[f'(x)=3\sin^2xcosx\]

OpenStudy (anonymous):

\[3\sin^2x=0\]

OpenStudy (anonymous):

\[cosx=0\]

OpenStudy (anonymous):

so my critical points for intervals of increase and decrease are

OpenStudy (anonymous):

\[\pi/2,-\pi/2\]

zepdrix (zepdrix):

Where are those 2 values coming from? From the cosine part?

OpenStudy (anonymous):

the cosx=0 at pi/2

OpenStudy (anonymous):

then the 3sin^2x=0, = 1-cosx/2=0 = cos2x=1

OpenStudy (anonymous):

shoot now im not sure, somehow the book has the interval of 0<x<pi

zepdrix (zepdrix):

\[cosx=0 \rightarrow x=\frac{ \pi }{ 2 }\pm \pi (n)\] Hmm your first set of critical points looks good. I think the next one you want to simply solve like this. \[3\sin^2x=0\] Dividing both sides by 3, then taking the square root of both sides, to get to this: \[sinx=0\rightarrow x=0 \pm \pi(n)\]

OpenStudy (anonymous):

ok, so my intervals are inc(0,pi/2) dec(pi/2,pi)

OpenStudy (anonymous):

my local max is f(pi/2) = 1

OpenStudy (anonymous):

ok so for second derivative, is it safe to say that 3sin^2x is f and cosx is g? (using the product rule)

zepdrix (zepdrix):

mhm that seems fine c:

OpenStudy (anonymous):